1. Suppose that the intersection of the AD curve and the SRAS curve lies to the right of the LRAS curve. The Keynesian prescription to resolve this problem would be ___________________ since the economy is _______________________.

a.a decrease in government expenditures, in an inflationary gap

b.an increase in taxes, in an inflationary gap

c.a decrease in taxes, in a recessionary gap

d.to do nothing, believed to be self-regulating

e.a or b

2. Exhibit 11-1

Taxable Income

Income Taxes

$0 - $9,000

10% of taxable income

$9,001 - $52,000

$900 + 15% of everything over $9,000

$52,001 - $300,000

$7,350 + 20% of everything over $52,000

Greater than $300,000

$56,950 + 30% of everything over $300,000


Refer to exhibit 11-1. The tax structure depicted in this tax schedule is a ______________ tax structure.

a.regressive

b.progressive

c.proportional

d.flat

3.

If government expenditures are $2,400, taxes are a flat 20 percent of Real GDP, current Real GDP is $7,200, and Natural Real GDP is $7,800. The cyclical deficit is _________________ and the structural deficit is _________________.

a.$120, $840

b.$960, $840

c.$840, $120

d.$840, $960

4.

Suppose that the economy is in a recessionary gap and the government increases its spending in an attempt to make the economy healthier. If there is complete crowding out it follows that

a.there will be no (net) shift of the AD curve.

b.the AD curve will shift rightward.

c.the AD curve will shift leftward.

d.the LRAS curve will shift rightward.

5. Supply-side economists assert that a decrease in marginal tax rates can shift the ___________ curve _______________, because lower marginal tax rates give people an incentive to __________________.

a.AD, rightward, purchase more consumer goods

b.AD, leftward, purchase fewer consumer goods

c.SRAS, leftward, engage in productive activities rather than leisure and tax-avoidance activities

d.SRAS, rightward, engage in productive activities rather than leisure and tax-avoidance activities

6. Exhibit 11-2

(Average)Tax Rate

Tax Base

Tax Revenues

40%

$200 billion

(A)

50%

(B)

$90 billion

60%

$140 billion

(C)


Refer to exhibit 11-2. What dollar amounts go in blanks (A), (B) and (C), respectively?

a.$120 billion, $45 billion, $56 billion

b.$500 billion, $180 billion, $233 billion

c.$240 billion, $40 billion, $200 billion

d.$80 billion, $180 billion, $84 billion

7. The budget deficit is a bigger percentage of GDP in year 1 than in year 2, so ________________.

a.the actual dollar budget deficit is bigger in year 1 than in year 2.

b.the actual dollar budget deficit is smaller in year 1 than in year 2.

c.the actual dollar budget deficit could be bigger or smaller in year 1 than in year 2.

d.the actual dollar budget deficit is the same in year 1 as in year 2.

please help me answer this right, most of my questions i posted here the answer been wrong.
Answer #1

1.

A

If short run equilibrium is in the right side of the LRAS, then it is inflationary gap in the economy. It will be resolved by decreasing the government spending. It will push the real GDP back to the long run potential GDP level.

2.

B

Tax rate is increasing with the increase in taxable income, then it is progressive tax rate system.

3.

A

4.

A

Increase in spending is offset by decrease in consumption and investment. It will cause no change in AD curve.

5.

D

Supply side economists favor policies for producers. When marginal tax applied to the producers, decreases, then SRAS shifts to the right and economy comes out of the recession as it also creates new jobs.

6.

D

Working note:

A = 40%*200 = $80 billion

B = 90/50% = $180 Billion

C = 140*60% = $84 Billion

Pl. repost other unanswered questions for their proper answers!